one car travels 300 miles while a car going ten miles faster travels 350. How fast is the first car going?
Answer by checkley77(12844)

D=RT
300=RT
350=(R+10)T
T=300/R
350=(R+10)300/R
350=(300R+3,000)/R
350R=300R+3,000
350R-300R=3,000
50R=3,000
R=3,000/50
R=60 MPH. FOR THE SLOWER CAR.
60+10=70 MPH FOR THE FASTER CAR.